# GS-ProCams: Gaussian Splatting-based Projector-Camera Systems

### v1

@@ 4.3 Ablation study
We conducted ablation study on the real-world benchmark dataset to validate the effectiveness of the projector’s PSF estimation (Eq. 8). We denote GS-ProCams without PSF as w/o PSF. The quantitative and qualitative results are shown in Tab. 3 and Fig. 8, respectively. Clearly, without estimating the projector’s PSF, the synthesized results ignores projector defocus, and exhibit unrealistically sharp projections, while GS-ProCams results are much closer to the camera-captured ground truth.
@@ 5 Conclusion
We introduce GS-ProCams, the first Gaussian Splatting-based framework for projector-camera systems (ProCams). This framework enables efficient novel viewpoint ProCams simulation and projector compensation without retraining. Compared with NeRF-based ProCams, our GS-ProCams waives the need for additional devices, and achieves higher ProCams simulation quality. Moreover, it is 600 times faster and uses only 1/10 of the GPU memory.
Limitations and future work. GS-ProCams has limitations in regions with strong specular reflections. More intricate photometric mapping calculations and optimization strategies are expected to further improve the framework’s performance, how to balance the simulation quality and efficiency is worth exploring in future work.
Supplementary Material
In the supplementary material, we provide a detailed description of the real-world data acquisition process in Appendix A, further implementation details of GS-ProCams in Appendix B, and additional qualitative results from experiments on ProCams simulation and projector compensation in Appendix C.
Appendix A Real-world data acquisition
We use an RGB camera to acquire data in indoor lighting environments. Since a fixed projector typically illuminates a specific region of the target, we randomly position multiple camera viewpoints along the hemispherical surface surrounding the target and capture data sequentially.
We employ the pinhole camera model to represent both the projector and the camera viewpoints. A digital speckle pattern as illustrated in Fig. 9 is projected for each viewpoint, enabling us to obtain the locations and rotations of the projector and the camera viewpoints using COLMAP [58]. When applying the trained GS-ProCams to a previously unseen viewpoint, we only need to capture an image from this viewpoint with the same projection and then register it into the existing COLMAP model.
Appendix B Implementation details
B.1 BRDF modeling
@@
B.2 Applications
As the principal manuscript outlines, our approach constitutes an efficient and fully differentiable framework for ProCams applications. In this section, we provide further details regarding the implementation.
Projector compensation. Our method facilitates view-agnostic projector compensation, as demonstrated in our experiments. Once the ProCams simulation is completed, GS-ProCams represents projector compensation as an inverse rendering problem. Given the desired appearance in the camera frame as $\mathbf{I}_{d}$, our framework allows for optimizing a virtual projector input pattern $\hat{\mathbf{I}}_{p}$ to make the simulation result $\hat{\mathbf{I}}_{c}$ in the virtual space align with $\mathbf{I}_{d}$ as following:
where $\hat{\mathbf{I}}^{*}_{p}$ is the compensated projector input for the desired visual display effect.
More applications. We present some additional applications of GS-ProCams as shown in Fig. 1, such as projection-based diminished reality and text-driven projection mapping, which are achieved by integrating GS-ProCams with popular large generative models. Specifically, for projection-based diminished reality, we eliminate objects captured in the camera frame using an inpainting model [65] and then apply the techniques as described in Sec. B.2 to obtain the projector input pattern, which can act on reality to remove the target objects visually. Similarly, we capture the base appearance of a target from a specific viewpoint, and then we can use text prompts to drive diffusion models [53, 21] for text-driven projection mapping. It is precisely because GS-ProCams efficiently models view-agnostic geometric and photometric mappings of ProCams that it can facilitate the application of ProCams in reality.
@@ B.2 Applications
Figure 10: Additional visual comparisons of ProCams simulation on the Nepmap synthetic dataset. The first column displays an object from a novel viewpoint, the second column shows the object under a novel projection pattern, and the third and fourth columns present the results of two methods.
Figure 11: Additional visual comparisons of projector compensation on real-world benchmark dataset. Each row illustrates the results of a scene under a novel viewpoint that our proposed GS-ProCams has not previously seen. The other two view-specific methods necessitate repeated data capturing and training for these viewpoints. The first three columns respectively illustrate the surface appearance from the specified viewpoint, the uncompensated appearance, and the expected appearance. The last three columns display the compensated results from different models.
Appendix C More results
We present additional qualitative results of ProCams simulation and projector compensation with novel projections under novel viewpoints that GS-ProCams has not seen previously. Additional visual comparisons of ProCams simulation on the Nepmap synthetic dataset [14] are shown in Fig. 11. Additional visual comparisons of ProCams simulation on real-world benchmark dataset are given in Fig. 12. Additional comparisons of projector compensation on real-world benchmark dataset are provided in Fig. 11. Please note that the presented results are based on a novel viewpoint and a novel projection that GS-ProCams has not seen before within each trained scene.
@@ Appendix C More results
Figure 12: Additional visual comparisons of ProCams simulation on real-world benchmark dataset. Each row illustrates the results of a target surface under a novel viewpoint that our proposed GS-ProCams has not previously seen, while DeProCams has retrained for it.

### v2

@@ 3.3 Ablation study
Simple projector PSF estimation. We denote the degraded version of GS-ProCams without simple PSF modeling (Eq.˜8) as GS-ProCams w/o PSF. In ProCams simulation, we observe that omitting PSF leads to noticeable artifacts on surfaces with complex geometry, as illustrated in Fig.˜8, and quantitative results are presented in Table˜3. For projector compensation, although the absence of PSF does not result in complete failure, it causes noticeable blurring in the results. The quantitative and qualitative results are shown in Table˜4 and Fig.˜7, respectively. This suggests the importance of PSF modeling in ProCams simulation, as discussed in [30, 36]. While we adopt a simple approximation, better simulation quality may be achieved by considering the insights in these works.
Number of training images. We evaluated the performance of GS-ProCams under varying numbers of textured training patterns on ProCams simulation. The number of training viewpoints was fixed at 25, and we varied the total number of training images by reducing the number of patterns per viewpoint. Note that GS-ProCams also includes a pure black projection at each training viewpoint, whereas our method uses joint optimization rather than multi-stage training as in Nepmap [13]. Quantitative results are presented in Table˜3. The results demonstrate that our method consistently outperforms view-specific baselines across different training pattern numbers. In contrast to the CNN-based method DeProCams [23], our method and DPCS [38] are both built on explicit modeling of ProCams, and do not require large amounts of training data to achieve high-quality ProCams simulation.
Number of training viewpoints. We further evaluated the performance of GS-ProCams under a reduced number of training viewpoints (fewer than 25) for ProCams simulation on novel viewpoints. To maintain a total of 100 textured training patterns, we increased the number of training patterns per viewpoint as the number of viewpoints decreased. Quantitative results are presented in Table˜5. The results demonstrate that GS-ProCams can still achieve high-quality ProCams simulation on unseen viewpoints, even when trained with as few as 4 viewpoints.
4 Limitations and Future Work
While our GS-ProCams framework demonstrates state-of-the-art performance in view-agnostic ProCams simulation and projector compensation, it also achieves significantly improved computational efficiency, as demonstrated in our exhaustive experiments (Section˜3). Nonetheless, several limitations remain. In this section, we discuss key challenges and possible directions for future work.
ProCams setup. Our method is designed for static projector-surface configurations and is not intended for dynamic scenes involving object motion or significant environmental lighting changes. In this setting, we jointly estimate the geometry and material of the projection surface for efficient view-agnostic projection mapping. As a result, regions outside the projector’s field of view, such as shadowed areas, may exhibit undesired artifacts, as accurate estimation in those regions is ill-posed under joint optimization. Moreover, our self-calibration process is based on SfM [60], which requires maximizing the camera’s coverage of the projection area.
Simplified model. Real-world scenes are inherently more complex, often involving strong global illumination effects such as inter-reflections and subsurface scattering. We address this in part by jointly estimating a view-dependent residual color term to account for global effects (Eq.˜11), but this remains insufficient. Our method cannot handle highly reflective or transparent materials, such as glass or strongly specular surfaces (Fig.˜9). In addition, the complex characteristics of the hardware system pose further challenges. For example, we estimate a very simple PSF term (Eq.˜8) for defocus blur, as validated in our ablation study (Fig.˜8). However, for objects with large depth variations, the projector’s shallow depth of field (DOF) introduces strong spatially-varying blur that GS-ProCams cannot fully model. Incorporating more accurate PSF estimation approaches [30, 36] with the GS-ProCams pipeline may better address this issue. Another point to note is that we find these factors may lead to ambiguities in both geometry and material parameters estimation, although GS-ProCams still achieves satisfactory ProCams simulation results.
Future work. To address the above limitations, future work could explore dynamic, view-agnostic ProCams that adapt to object or projector motion as well as lighting changes. Additionally, more precise optical and lighting models may further enhance realism and accuracy.
Figure 9: Failure case. GS-ProCams is not capable of handling complex materials such as transparent glass and highly specular surfaces, which exhibit strong view-dependent transmittance and reflections.
@@ 5 Conclusion
We introduce GS-ProCams, the first Gaussian Splatting-based framework for projector-camera systems (ProCams). This framework enables efficient view-agnostic ProCams simulation and projector compensation without retraining. Compared with NeRF-based ProCams, our GS-ProCams simplifies the requirements and supports environmental illumination. While GS-ProCams achieves higher ProCams simulation quality. Moreover, it is 900 times faster and uses only 1/10 of the GPU memory.
Appendix A Introduction
In the supplementary material, we provide a detailed description of the real-world data acquisition process in Appendix˜B, further implementation details of GS-ProCams in Appendix˜C, and additional results from experiments on ProCams simulation and projector compensation in Appendix˜D.
Appendix B Real-world data acquisition
We use an RGB camera to acquire data in indoor lighting environments. Since a fixed projector illuminates a specific region, we randomly position multiple camera viewpoints along the hemispherical surface surrounding the projection surface and capture data sequentially.
We employ the pinhole camera model to represent both the projector and the camera viewpoints. A digital speckle pattern as illustrated in Fig.˜10 is projected for each viewpoint, enabling us to obtain the locations and rotations of the projector and the camera viewpoints using COLMAP [60]. When applying the trained GS-ProCams to a previously unseen viewpoint, we only need to capture an image from this viewpoint with the same projection and then register it into the existing COLMAP model.
Appendix C Implementation details
C.1 BRDF modeling
@@
C.2 Applications
As the principal manuscript outlines, our approach constitutes an efficient and fully differentiable framework for ProCams applications. Our compensation framework serves as the foundation for diverse projection mapping applications.
We present some additional applications of GS-ProCams as shown in the principal manuscript, such as projection-based diminished reality and text-driven projection mapping, which are achieved by integrating GS-ProCams with popular large generative models. Specifically, for projector-based diminished reality, we eliminate objects captured in the camera frame using an inpainting model [65] and then apply our compensation method to obtain the projector input pattern, which can act on reality to remove the target objects visually. Similarly, we capture the base appearance of a target from a specific viewpoint, and then we can use text prompts to drive diffusion models [55, 20] for text-driven projection mapping. It is precisely because GS-ProCams efficiently models view-agnostic geometric and photometric mappings of ProCams that it can facilitate the application of ProCams in reality.
Appendix D More results
Procams simulation. We present the quantitative per-scene results on the Nepmap synthetic dataset [13] in Table˜6. Additional visual comparisons of ProCams simulation on the synthetic dataset are shown in Fig.˜11. Additional visual comparisons of ProCams simulation on real-world benchmark dataset are given in Fig.˜12.
Projector compensation. To further evaluate the reliability of GS-ProCams compensation under novel viewpoints, we test it on the same scene using the same projector input across different camera views, as illustrated in Fig.˜14. Additional comparisons of projector compensation on different surfaces are provided in Fig.˜14.
@@ Appendix D More results
Table 6: ProCams simulation results on the Nepmap synthetic dataset [13]. Since synthetic scenes ignore defocus blur, we use a degraded version of GS-ProCams that excludes the projector PSF for these experiments.
Zoo Castle Pear Planck Average
[…]
Nepmap [13] 25.68 / 0.9300 / 0.046 28.93 / 0.9404 / 0.065 27.29 / 0.9648 / 0.062 27.06/ 0.9427 / 0.071 $\sim$3 hrs >>24 GB 0.3
GS-ProCams w/o PSF 27.68 / 0.9529 / 0.043 35.42 / 0.9772 / 0.031 34.00 / 0.9803 / 0.031 31.40 / 0.9676 / 0.041 3.9 min 2.03 GB 275
Adopted from the original paper
@@
Figure 11: Additional visual comparisons of ProCams simulation on the Nepmap synthetic dataset [13]. The first column displays an object from a novel viewpoint without projection, the second column shows the object under a novel projection pattern, and the third and fourth columns present the simulated results from two methods.
@@
Figure 12: Additional visual comparisons of ProCams simulation on real-world benchmark dataset. The top row shows results from a camera viewpoint used during GS-ProCams training, while the bottom row shows results from a novel viewpoint that GS-ProCams has not seen during training. The first column shows the projector input pattern, followed by the projection surface under uniform gray projection. The remaining columns present the ground truth and simulation results from each method.
Figure 13: Visual comparisons of projector compensation across novel viewpoints. Each row presents a different viewpoint of the same surface. The first three columns show the projection surface, the uncompensated appearance, and the desired appearance. The remaining columns display the camera-captured compensated results for the same projection from different methods. While view-specific methods require additional data acquisition and retraining for each novel viewpoint, GS-ProCams generalizes effectively, delivering consistent and high-quality results across unseen views.
Figure 14: Additional visual comparisons of projector compensation. Each row illustrates the results of a scene under a novel viewpoint that our proposed GS-ProCams has not previously seen. The other three view-specific methods necessitate repeated data capturing and training for these viewpoints. The first three columns respectively illustrate the projection surface from the specified viewpoint, the uncompensated appearance, and the expected appearance. The remaining columns display the compensated results from different models.